Consider f; Z+ → Z+ define by f(a) =a2. Check if f is one-to-one and / or into using suitable explanation


1
Expert's answer
2021-09-09T15:06:07-0400

One to one functions are special functions that return a unique range for each element in their domain, implying that the responses never duplicate.

"f(a)=a^2\\
Consider \space F(a)=F(b)\\
a^2=b^2\\
a=b\\"

F is one-to-one